A married couple are trying to finance their four-year-old son's college education. Money can be deposited at 9 % compounded quarterly. What end-of-quarter deposit must be made from the son's 4th birthday to his 18th birthday to provide $60,000 on each birthday from the 18th to the 21st? (Note that the first deposit comes three months after his 4th birthday and the last deposit is made on the date of the first withdrawal.)
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

B Annual interest rate Quarterly interest rate Effective Annual rate 9% 2.25% 9.31% - moro $60,000.00 Annual withdrawal from

Cell reference -

в Annual interest rate Quarterly interest rate Effective Annual rate 0.09 =C2/4 =(1+C3)^4-1 60000 Annual withdrawal from 18th

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.